Solve each system by substitution. Check you solution
0.5x + 0.25y = 36
y + 18 = 16x

Answer by mukhopadhyay(490)

y + 18 = 16x
=>y=16x-18
0.5x + 0.25y = 36 (multiply both sides by 4)
=>2x+y = 144
Substitute y=16x-8
=>2x+16x-8=144
=>18x=152
=>x=152/18=9
y=16x-18
=>y=16(9)-18
=>y=144-18=126
Answer: x=9 and y=126